OpenStudy (anonymous):

Evaluate 3{5 + 3[10 + 4 · 8]}

OpenStudy (anonymous):

First : [10 + 4 · 8]=? And in there first 4 * 8 K ? What is the answer there ?;) @Jmat1212

OpenStudy (anonymous):

112

OpenStudy (anonymous):

32

OpenStudy (anonymous):

42

OpenStudy (anonymous):

WE have : [10 +( 4 · 8 )]=10+32 = 42 Then : 3*(42)=126 then : 126+5=131 then : 3(131)=393 Got it ?;) You had a small fall in your answer !;)

OpenStudy (anonymous):

i got it now thank you;)

OpenStudy (anonymous):

one more ?